# Sharp asymptotics for $N$-point correlation functions of coalescing heavy-tailed random walks

### v1

@@ 3 First collision time of random walks
$\begin{array}[]{l}\displaystyle\quad\sum_{y_{ij}}p(y_{ij})\Big{|}{\mathbb{P}}(\tau_{\bf x}>s,\tau_{ij}>t,\Delta^{ij}_{t}=y_{ij})-\frac{p_{\rm NC}(s)}{B_{2t}\hskip 1.0ptl(2t)}\Big{|}\\[5.0pt] \displaystyle=\!\sum_{z_{ij}\in D_{1}}\!\!+\!\!\sum_{z_{ij}\in D_{2}}\!{\mathbb{P}}(\tau_{\bf x}>s,\Delta^{ij}_{s}=z_{ij})\sum_{y_{ij}}p(y_{ij})\Big{|}{\mathbb{P}}_{y_{ij}}(\hat{\tau}_{ij}>t-s,\widehat{\Delta}_{t-s}^{ij}=z_{ij})-\frac{1}{B_{2t}\hskip 1.0ptl(2t)}\Big{)}\Big{|},\end{array}$
where $D_{1}:=\{z:|z|>B_{s}\ell(t)^{\theta_{3}}\}$ and $D_{2}=\{z:|z|\leq B_{s}\ell(t)^{\theta_{3}}\}$ with $\theta_{3}>0$ to be chosen later, we decomposed at time $s$, and considered the time-reversed rate 2 random walk $\widehat{\Delta}^{ij}$ during $(s,t]$. By the large deviation and (2.34), the sum for $z_{ij}$ over $D_{1}$ is bounded above by
$\frac{C}{B_{t}\hskip 1.0ptl(t)}C\hskip 1.0pt{\mathbb{P}}(\Delta_{s}^{ij}\in D_{2})\leq\frac{Cs\hskip 1.0ptL\big{(}B_{s}\ell(t)^{\theta_{3}}\big{)}}{B_{t}\hskip 1.0ptl(t)B_{s}^{\alpha}\ell(t)^{\alpha\theta_{3}}}\leq\frac{C}{B_{t}\hskip 1.0ptl({t})\ell(t)^{\alpha\theta_{3}-\varepsilon}}.$ (3.15)
On $D_{2}$, we utilize Lemma 2.13 and then consider $\sum_{y_{ij}}p(y_{ij}){\mathbb{P}}_{y_{ij}}(\hat{\tau}_{ij}>t-s){\mathbb{P}}_{0}(\widehat{\Delta}^{ij}_{t-s}=0)$ instead. On the one hand, we use Lemma 2.9 to approximate $\sum_{y_{ij}}p(y_{ij}){\mathbb{P}}_{y_{ij}}(\hat{\tau}_{ij}>t-s)$ by $\frac{1}{g_{\alpha}(0)\hskip 1.0ptl(2t-2s)}$, where the difference of the last term and $\frac{1}{g_{\alpha}(0)\hskip 1.0ptl(2t)}$ is O($\frac{1}{l(t)\ell(t)}$) by (2.27). On the other hand, we replace ${\mathbb{P}}_{0}(\widehat{\Delta}^{ij}_{t-s}=0)$ by $g_{\alpha}(0)/B_{2t}$ with an error term $O((t/s)^{-\frac{1}{1+\alpha}+\varepsilon}/B_{t})$ thanks to (1.5) and Proposition 2.3. Combining all these bounds, we have that the sum for $z_{ij}$ over $D_{2}$ is bounded above by
$Cp_{\rm NC}(s)\Big{(}\frac{1}{B_{t}\hskip 1.0ptl(t)\ell(t)^{\theta_{4}}}+\frac{B_{s}\ell(t)^{\theta_{3}}}{B_{t}^{2}}+\frac{1}{B_{t}\hskip 1.0ptl(t)\ell(t)}+\frac{(t/s)^{-\frac{1}{1+\alpha}+\varepsilon}}{B_{t}\hskip 1.0ptl(t)}\Big{)},$ (3.16)
@@
$\frac{C}{B_{t}\hskip 1.0ptl(t)}\Big{(}\frac{p_{\rm NC}(s)}{\ell(t)}+\frac{1}{\ell(t)^{\kappa}}\Big{)}.$
@@
${\mathbb{P}}(\tau_{{\bf x}}>t)\leq\frac{C_{\bf x}}{l(t)^{\upsilon_{N}}}\quad\mbox{and}\quad\Big{|}\frac{\mathrm{d}}{\mathrm{d}t}{\mathbb{P}}(\tau_{{\bf x}}>t)\Big{|}\leq\frac{C_{\bf x}}{B_{t}\hskip 1.0ptl(t)^{\upsilon_{N}+1}}.$ (3.17)
Proof For $\alpha=1$, note that $l(t)\leq\ell(t)^{K}$ as $B_{t}\in{\cal B}$. We take $\kappa=K(\upsilon_{N}+1)$ in Lemma 3.1 and then insert (3.2) into the ODE (3.1). This yields that for any $\varepsilon>0$,
$-\frac{\mathrm{d}}{\mathrm{d}t}p_{\rm NC}(t)\geq\frac{2\upsilon_{N}-C\hskip 1.0pt\ell(t)^{-1+\varepsilon}}{B_{2t}l(2t)}p_{\rm NC}(s)-\frac{C}{B_{t}l(t)^{\upsilon_{N}+2}}.$ (3.18)
Define $F(t):=\int_{0}^{t}\frac{C}{B_{2s}\hskip 1.0ptl(2s)\ell(s)^{1-\varepsilon}}\mathrm{d}s$. Since
$\int_{0}^{t}\frac{\mathrm{d}s}{B_{2s}\hskip 1.0ptl(2s)\ell(s)^{1-\varepsilon}}\leq\frac{C\hskip 1.0pt\mathrm{d}l(s)}{\hskip 1.0ptl(s)^{1+(1-\varepsilon)/K}}<\infty,$
$F(t)$ is bounded. As $\big{(}\log(l(2t)^{\upsilon_{N}})\big{)}^{\prime}=\frac{2\upsilon_{N}}{B_{2t}l(2t)}$, multiplying to (3.18) the integrating factor $\exp\big{(}\log(l(2t)^{\upsilon_{N}})-F(t)\big{)}=e^{-F(t)}l(2t)^{\upsilon_{N}}$ and using the boundedness of $e^{-F(t)}$, we obtain
@@
The r.h.s. has a bounded integral, which in turn implies the first inequality in (3.17).
Similar calculations as in the proof of Lemma 2.10 yield
$|l(t)^{-\upsilon_{N}}-l(s)^{-\upsilon_{N}}|\leq\frac{C_{N}(t/s)^{\varepsilon}}{l(t)^{\upsilon_{N}}\ell(t)}.$ (3.19)
Now for $s=t/\ell(t)^{\theta}$ in Lemma 3.1, using the above bound and $p_{\rm NC}(s)\leq C\hskip 1.0ptl(2s)^{-\upsilon_{N}}$, and applying (3.2) into the ODE (3.1), we obtain
$\Big{|}\frac{\mathrm{d}}{\mathrm{d}t}{\mathbb{P}}(\tau_{{\bf x}}>t)\Big{|}\leq\frac{C_{\bf x}}{B_{t}\hskip 1.0ptl(t)^{\upsilon_{N}+1}}.$ (3.20)
@@
Proof of Theorem 1.4 Take $s=t/\ell(t)^{\theta}$ as in Lemma 3.1. Then, by Lemma 3.2 and (3.19),
${\mathbb{P}}(\tau_{{\bf x}}\in(s,t])\leq\int_{s}^{t}\frac{C_{\bf x}}{B_{u}\hskip 1.0ptl(u)^{\upsilon_{N}+1}}\mathrm{d}u\leq\frac{C_{\bf x}}{l(t)^{\upsilon_{N}}\ell(t)^{1-\varepsilon}}.$ (3.21)
Therefore, by (3.1) and (3.2), there exists some $\delta>0$ such that
$-\frac{\mathrm{d}}{\mathrm{d}t}{\mathbb{P}}(\tau_{{\bf x}}>t)=\frac{2\upsilon_{N}}{B_{2t}l(2t)}{\mathbb{P}}(\tau_{{\bf x}}>t)+R(t),$
@@
[…]
@@
${\mathbb{P}}_{\bf x}(\tau_{{\bf x}}>t)=\frac{c_{\bf x}}{l(2t)^{\upsilon_{N}}}\Big{(}1+O({\cal R}(t))\Big{)},$ (3.22)
where ${\cal R}(t)=\int_{t}^{\infty}\frac{\mathrm{d}u}{u\hskip 1.0pt\ell(u)^{2-\varepsilon}}=o\big{(}\ell(t)^{-2+\varepsilon}\big{)}$.
For $\alpha<1$, recall that $B_{t}$ is $1/\alpha$-regularly varying. The same argument implies that (3.22) holds with ${\cal R}(t)=t^{-\eta}$, where $\eta$ is a constant in $(1/\alpha-1,\hskip 1.0pt\delta+1/\alpha-1)$.
@@ 3.2 Asymptotic independence
Under the assumptions of Theorem 1.4, for independent random walks $({\mathbf{X}}_{t})_{t\geq 0}$, there exist $C_{\bf x}$ and $\delta>0$ such that for any $\varepsilon>0$,
$\sum_{{\bf y}\in{\mathbb{Z}}^{N}}\Big{|}{\mathbb{P}}_{\bf x}({\mathbf{X}}_{t}={\bf y},\tau_{{\bf x}}>t)-{\mathbb{P}}_{\bf x}({\mathbf{X}}_{t}={\bf y}){\mathbb{P}}_{\bf x}(\tau_{{\bf x}}>t)\Big{|}\leq\left\{\begin{array}[]{ll}\displaystyle\frac{C_{\bf x}}{l(t)^{\upsilon_{N}}\hskip 1.0pt\ell(t)^{1-\varepsilon}},&\alpha=1;\\[10.0pt] \displaystyle\frac{C_{\bf x}}{l(t)^{\upsilon_{N}}\hskip 1.0ptt^{\delta}},&\alpha\in(0,1).\end{array}\right.$ (3.23)
Remark 3.4
Via a simple adaption of the proof below, we can change the fixed ${\bf x}$ to the origin ${\bf 0}$ such that the same upper bound as in (3.23) holds for
$\sum_{{\bf y}\in{\mathbb{Z}}^{N}}\Big{|}{\mathbb{P}}_{\bf x}({\mathbf{X}}_{t}={\bf y},\tau_{{\bf x}}>t)-{\mathbb{P}}_{\bf 0}({\mathbf{X}}_{t}={\bf y}){\mathbb{P}}_{\bf x}(\tau_{{\bf x}}>t)\Big{|}$
Proof Let $s=t/\ell(t)^{\theta_{1}}$ with $\theta_{1}<\frac{\alpha}{1-\alpha}$. For the l.h.s. of (3.23), we bound the summand for each ${\bf y}$ by the sum of the following four differences.
@@
The same calculations as in (3.21) lead to that for any $\varepsilon>0$,
$\sum_{{\bf y}}(R^{\bf y}_{1}+R^{\bf y}_{2})={\mathbb{P}}(\tau_{{\bf x}}\in(s,t])\leq\frac{C_{\bf x}}{l(t)^{\upsilon_{N}}\ell(t)^{1-\varepsilon}}.$ (3.24)
@@
$\sum_{{\bf y}}R_{3}^{\bf y}\leq\sum_{{\bf y}}\sum_{\bf z}{\mathbb{P}}_{\bf x}({\mathbf{X}}_{s}={\bf z},\tau_{{\bf x}}>s)\big{|}{\mathbb{P}}_{{\bf z}}({\mathbf{X}}_{t-s}={\bf y})-{\mathbb{P}}_{{\bf x}}({\mathbf{X}}_{t-s}={\bf y})\big{|}.$ (3.25)
A similar bound also holds for $R_{4}^{\bf y}$. We then bound the r.h.s. of (3.25).
For ${\bf x}=(x_{1},\ldots,x_{N})$, denote $|{\bf x}|_{\rm m}:=\max_{1\leq i\leq N}|x_{i}|$. Let $\theta_{2}>0$. Let $D_{1}:=\{\textbf{w}\in{\mathbb{Z}}^{d}:|\textbf{w}-{\bf x}|_{\rm m}\leq B_{t}\hskip 1.0pt\ell(t)^{\theta_{2}}\}$ and $D_{2}:=\{\textbf{w}\in{\mathbb{Z}}^{d}:|\textbf{w}-{\bf x}|_{\rm m}\leq B_{s}\hskip 1.0pt\ell(t)^{\theta_{2}}\}$. We consider three regions: (a) $\{{\mathbf{X}}_{t}\in D_{1}^{c}\}$, (b) $\{{\mathbf{X}}_{s}\in D_{2}^{c}\}$, and (c) $\{{\mathbf{X}}_{t}\in D_{1},{\mathbf{X}}_{s}\in D_{2}\}$. The large deviation (2.10) and basic estimates, such as the Potter bound and (1.6), give a bound of the probabilities of the first two events (a) and (b):
${\mathbb{P}}({\mathbf{X}}_{t}\in D_{1}^{c})+{\mathbb{P}}({\mathbf{X}}_{s}\in D_{2}^{c})\leq\frac{C_{\bf x}}{\ell(t)^{\alpha\theta_{2}-\varepsilon}}.$ (3.26)
@@
$\begin{array}[]{l}\displaystyle\sum_{{\bf y}\in D_{1}}\sum_{{\bf z}\in D_{2}}{\mathbb{P}}_{\bf x}({\mathbf{X}}_{s}={\bf z},\tau_{{\bf x}}>s)\big{|}{\mathbb{P}}_{{\bf z}}({\mathbf{X}}_{t-s}={\bf y})-{\mathbb{P}}_{{\bf x}}({\mathbf{X}}_{t-s}={\bf y})\big{|}\\[5.0pt] \displaystyle\quad\leq\frac{C_{\bf x}B_{s}\hskip 1.0pt\ell(t)^{\theta_{2}}}{B_{t}^{N+1}}\sum_{{\bf y}\in D_{1}}\sum_{{\bf z}\in D_{2}}{\mathbb{P}}_{\bf x}({\mathbf{X}}_{s}={\bf z},\tau_{{\bf x}}>s)\leq\frac{C_{\bf x}B_{s}B_{t}^{N}\hskip 1.0pt\ell(t)^{(N+1)\theta_{2}}}{B_{t}^{N+1}\hskip 1.0ptl(t)^{\upsilon_{N}}},\end{array}$ (3.27)
where in the last inequality we used that the sum over ${\bf z}$ is bounded by ${\mathbb{P}}(\tau_{x}>s)\leq C_{\bf x}l(t)^{-\upsilon_{N}}$. For $\alpha=1$, if we take $\theta_{2}=K\upsilon_{N}+1$ and $\theta_{1}=(N+1)\theta_{2}+1$, then combining the three bounds (3.24), (3.26) and (3.27) we arrive at the desired bound (3.23). For $\alpha<1$, we take $\theta_{2}$ sufficiently small and $\theta_{1}=(N+1)\alpha\theta_{2}+\alpha$ such that $\theta_{1}<\frac{\alpha}{1-\alpha}$ is satisfied. Then, (3.24), (3.26) and (3.27) imply that there exists some $\delta$ such that (3.23) holds, which finishes the proof.

### v2

@@ 3 First collision time of random walks
$\begin{array}[]{l}\displaystyle\quad\Big|\sum_{y_{ij}}p(y_{ij}){\mathbb{P}}(\tau_{\bf x}>s,\tau_{ij}>t,\Delta^{ij}_{t}=y_{ij})-\frac{p_{\rm NC}(s)}{B_{2t}\hskip 1.0ptl(2t)}\Big|\\[5.0pt] \displaystyle=\!\sum_{z_{ij}\in D_{1}}\!\!+\!\!\sum_{z_{ij}\in D_{2}}\!{\mathbb{P}}(\tau_{\bf x}>s,\Delta^{ij}_{s}=z_{ij})\Big|\sum_{y_{ij}}p(y_{ij}){\mathbb{P}}_{y_{ij}}(\hat{\tau}_{ij}>t-s,\widehat{\Delta}_{t-s}^{ij}=z_{ij})-\frac{1}{B_{2t}\hskip 1.0ptl(2t)}\Big|,\end{array}$
where $D_{1}:=\{z:|z|>B_{s}\ell(t)^{\theta_{3}}\}$ and $D_{2}=\{z:|z|\leq B_{s}\ell(t)^{\theta_{3}}\}$ with $\theta_{3}>0$ to be chosen later. To obtain the above, we have decomposed at time $s$, and considered the time-reversed rate 2 random walk $\widehat{\Delta}^{ij}$ during $(s,t]$. By using (56) and then large deviations, the sum for $z_{ij}$ over $D_{1}$ is bounded above by
$\frac{C}{B_{t}\hskip 1.0ptl(t)}\hskip 1.0pt{\mathbb{P}}(\Delta_{s}^{ij}\in D_{2})\leq\frac{Cs\hskip 1.0ptL\big(B_{s}\ell(t)^{\theta_{3}}\big)}{B_{t}\hskip 1.0ptl(t)B_{s}^{\alpha}\ell(t)^{\alpha\theta_{3}}}\leq\frac{C}{B_{t}\hskip 1.0ptl({t})\ell(t)^{\alpha\theta_{3}-\varepsilon}}.$ (78)
On $D_{2}$, firstly, by (57) in Lemma 2.13 and by the fact that $|z_{ij}|\leq B_{s}\ell(t)^{\theta_{3}}$,
$\sum_{y_{ij}}p(y_{ij})\big|{\mathbb{P}}_{y_{ij}}(\hat{\tau}_{ij}>t-s,\widehat{\Delta}_{t-s}^{ij}=z_{ij})-{\mathbb{P}}_{y_{ij}}(\hat{\tau}_{ij}>t-s){\mathbb{P}}_{0}(\widehat{\Delta}_{t-s}^{ij}=0)\big|\leq\frac{C}{B_{t}\hskip 1.0ptl(t)\ell(t)^{\theta_{4}}}+\frac{C\hskip 1.0ptB_{s}\ell(t)^{\theta_{3}}}{B_{t}^{2}},$
where we can choose any $\theta_{4}<\alpha/(1-\alpha)$. Secondly, by Lemma 2.9 and recalling from (5) the expression ${\mathbb{P}}_{0}(\widehat{\Delta}_{t-s}^{ij}=0)=g_{\alpha}(0)/B_{2t-2s}$, we have
$\Big|\sum_{y_{ij}}p(y_{ij}){\mathbb{P}}_{y_{ij}}(\hat{\tau}_{ij}>t-s){\mathbb{P}}_{0}(\widehat{\Delta}_{t-s}^{ij}=0)-\frac{1}{B_{2t}l(2t)}\Big|=\Big|\frac{1+O(1/\ell(t))}{g_{\alpha}(0)\hskip 1.0ptl(2t-2s)}\frac{g_{\alpha}(0)}{B_{2t-2s}}-\frac{1}{B_{2t}l(2t)}\Big|.$
We then insert $\frac{1}{l(2t-2s)}=\frac{1}{l(2t)}+O(\frac{1}{l(t)\ell(t)})$ by (46) and insert $\frac{1}{B_{2t-2s}}=\frac{1}{B_{2t}}+O(\frac{s}{tB_{t}})$ by Proposition 2.3 in the above equality. This, together with the error bound in the first step, yields that the sum for $z_{ij}$ over $D_{2}$ is bounded above by
$C\hskip 1.0ptp_{\rm NC}(s)\Big(\frac{1}{B_{t}\hskip 1.0ptl(t)\ell(t)^{\theta_{4}}}+\frac{B_{s}\ell(t)^{\theta_{3}}}{B_{t}^{2}}+\frac{1}{B_{t}\hskip 1.0ptl(t)\ell(t)}+\frac{(s/t)}{B_{t}\hskip 1.0ptl(t)}\Big)$ (79)
@@
$\frac{C}{B_{t}\hskip 1.0ptl(t)}\Big(\frac{p_{\rm NC}(s)}{\ell(t)^{1-\varepsilon}}+\frac{1}{\ell(t)^{\kappa}}\Big).$
@@
${\mathbb{P}}(\tau_{{\bf x}}>t)\leq\frac{C_{\bf x}}{l(t)^{\upsilon_{N}}}\quad\mbox{and}\quad\Big|\frac{\mathrm{d}}{\mathrm{d}t}{\mathbb{P}}(\tau_{{\bf x}}>t)\Big|\leq\frac{C_{\bf x}}{B_{t}\hskip 1.0ptl(t)^{\upsilon_{N}+1}}.$ (80)
Proof For $\alpha=1$, note that $l(t)\leq\ell(t)^{K}$ as $B_{t}\in{\cal B}$. We take $\kappa=K(\upsilon_{N}+1)$ in Lemma 3.1 and then insert (64) into the ODE (63). This yields that for any $\varepsilon>0$,
$\frac{\mathrm{d}}{\mathrm{d}t}p_{\rm NC}(t)\geq\frac{2\upsilon_{N}-C\hskip 1.0pt\ell(t)^{-1+\varepsilon}}{B_{2t}l(2t)}p_{\rm NC}(t)-\frac{C}{B_{t}l(t)^{\upsilon_{N}+2}},$ (81)
where we used $p_{\rm NC}(s)\geq p_{\rm NC}(t)$. Define $F(t):=\int_{0}^{t}\frac{C}{B_{2s}\hskip 1.0ptl(2s)\ell(s)^{1-\varepsilon}}\mathrm{d}s$, which is bounded since
$\int_{0}^{t}\frac{\mathrm{d}s}{B_{2s}\hskip 1.0ptl(2s)\ell(s)^{1-\varepsilon}}\leq\int_{0}^{\infty}\frac{C\hskip 1.0pt\mathrm{d}l(s)}{\hskip 1.0ptl(s)^{1+(1-\varepsilon)/K}}<\infty.$
As $\big(\log(l(2t)^{\upsilon_{N}})\big)^{\prime}=\frac{2\upsilon_{N}}{B_{2t}l(2t)}$, multiplying to (81) the integrating factor $\exp\big(\log(l(2t)^{\upsilon_{N}})-F(t)\big)=e^{-F(t)}l(2t)^{\upsilon_{N}}$ and using $e^{-F(t)}\leq 1$, we obtain
@@
The r.h.s. has a bounded integral. This and boundedness of $F(t)$ imply the first inequality in (80).
To prove the second inequality in (80), we need to derive a bound in the opposite direction to that of (80) for the derivative of $p_{\rm NC}(t)$. Similar calculations as in the proof of Lemma 2.10 yield
$|l(t)^{-\upsilon_{N}}-l(s)^{-\upsilon_{N}}|\leq\frac{C_{N}(t/s)^{\varepsilon}}{l(t)^{\upsilon_{N}}\ell(t)}.$ (82)
Now for $s=t/\ell(t)^{\theta}$ in Lemma 3.1, using the above bound and $p_{\rm NC}(s)\leq C\hskip 1.0ptl(2s)^{-\upsilon_{N}}$, and applying (64) into the ODE (63), we obtain
$\Big|\frac{\mathrm{d}}{\mathrm{d}t}{\mathbb{P}}(\tau_{{\bf x}}>t)\Big|\leq\frac{C_{\bf x}}{B_{t}\hskip 1.0ptl(t)^{\upsilon_{N}+1}}.$ (83)
@@
Proof of Theorem 1.2 Take $s=t/\ell(t)^{\theta}$ as in Lemma 3.1. Then, by Lemma 3.2 and (82),
${\mathbb{P}}(\tau_{{\bf x}}\in(s,t])\leq\int_{s}^{t}\frac{C_{\bf x}}{B_{u}\hskip 1.0ptl(u)^{\upsilon_{N}+1}}\mathrm{d}u\leq\frac{C_{\bf x}}{l(t)^{\upsilon_{N}}\ell(t)^{1-\varepsilon}}.$ (84)
Therefore, by (63) and (64), there exists some $\delta>0$ such that
$\frac{\mathrm{d}}{\mathrm{d}t}{\mathbb{P}}(\tau_{{\bf x}}>t)=\frac{2\upsilon_{N}}{B_{2t}l(2t)}{\mathbb{P}}(\tau_{{\bf x}}>t)+R(t),$
@@
For $\alpha=1$, since $\frac{1}{t\hskip 1.0pt\ell(t)^{2-\varepsilon}}=\frac{1}{B_{t}\hskip 1.0ptl(t)\ell(t)^{1-\varepsilon}}\leq\frac{1}{B_{t}\hskip 1.0ptl(t)^{1+\varepsilon/K}}$ is integrable, by Karamata’s theorem $\int_{t}^{\infty}\frac{\mathrm{d}u}{u\hskip 1.0pt\ell(u)^{2-\varepsilon}}\to 0$ is slowly varying and $\frac{1}{\ell(t)^{2-\varepsilon}}\int_{t}^{\infty}\frac{\mathrm{d}u}{u\hskip 1.0pt\ell(u)^{2-\varepsilon}}\to\infty$. We then have
@@
${\mathbb{P}}_{\bf x}(\tau_{{\bf x}}>t)=\frac{c_{\bf x}}{l(2t)^{\upsilon_{N}}}\Big(1+O({\cal R}_{\rm NC}({\bf x},t))\Big),$ (85)
where ${\cal R}_{\rm NC}({\bf x},t)=C_{\bf x}\int_{t}^{\infty}\frac{\mathrm{d}u}{u\hskip 1.0pt\ell(u)^{2-\varepsilon}}=o\big(\ell(t)^{-2+\varepsilon}\big)$.
For $\alpha<1$, recall that $B_{t}$ is $1/\alpha$-regularly varying. The same argument implies that (85) holds with ${\cal R}(t)=C_{\bf x}t^{-\eta}$, where $\eta$ is a constant in $(1/\alpha-1,\hskip 1.0pt\delta+1/\alpha-1)$.
@@ 3.2 Asymptotic independence
Under the assumptions of Theorem 1.2, for independent random walks $({\mathbf{X}}_{t})_{t\geq 0}$, there exist $C_{\bf x}$ and $\delta>0$ such that for any $\varepsilon>0$ and sufficiently large $t$,
$\sum_{{\bf y}\in{\mathbb{Z}}^{N}}\Big|{\mathbb{P}}_{\bf x}({\mathbf{X}}_{t}={\bf y},\tau_{{\bf x}}>t)-{\mathbb{P}}_{\bf x}({\mathbf{X}}_{t}={\bf y}){\mathbb{P}}_{\bf x}(\tau_{{\bf x}}>t)\Big|\leq\left\{\begin{array}[]{ll}\displaystyle\frac{C_{\bf x}}{l(t)^{\upsilon_{N}}\hskip 1.0pt\ell(t)^{1-\varepsilon}},&\alpha=1;\\[10.0pt] \displaystyle\frac{C_{\bf x}}{l(t)^{\upsilon_{N}}\hskip 1.0ptt^{\delta}},&\alpha\in(0,1).\end{array}\right$. (86)
Proof Let $s=t/\ell(t)^{\theta_{1}}$ with $\theta_{1}<\frac{\alpha}{1-\alpha}$. For the l.h.s. of (86), we bound the summand for each $\bf y$} by the sum of the following four differences.
@@
The same calculations as in (84) lead to the fact that for any $\varepsilon>0$,
$\sum_{{\bf y}\in{\mathbb{Z}}^{N}}(R^{\bf y}_{1}+R^{\bf y}_{2})=2{\mathbb{P}}_{\bf x}(\tau_{{\bf x}}\in(s,t])\leq\frac{C_{\bf x}}{l(t)^{\upsilon_{N}}\ell(t)^{1-\varepsilon}}.$ (87)
@@
$\sum_{{\bf y}}R_{3}^{\bf y}\leq\sum_{{\bf y}}\sum_{\bf z}{\mathbb{P}}_{\bf x}({\mathbf{X}}_{s}={\bf z},\tau_{{\bf x}}>s)\big|{\mathbb{P}}_{{\bf z}}({\mathbf{X}}_{t-s}={\bf y})-{\mathbb{P}}_{{\bf x}}({\mathbf{X}}_{t-s}={\bf y})\big|.$ (88)
A similar bound also holds for $R_{4}^{\bf y}$. We then bound the r.h.s. of (88).
For ${\bf x}=(x_{1},\ldots,x_{N})$, denote $|{\bf x}|_{\rm m}:=\max_{1\leq i\leq N}|x_{i}|$. Let $\theta_{2}>0$. Let $D_{1}:=\{\textbf{w}\in{\mathbb{Z}}^{d}:|\textbf{w}-{\bf x}|_{\rm m}\leq B_{t}\hskip 1.0pt\ell(t)^{\theta_{2}}\}$ and $D_{2}:=\{\textbf{w}\in{\mathbb{Z}}^{d}:|\textbf{w}-{\bf x}|_{\rm m}\leq B_{s}\hskip 1.0pt\ell(t)^{\theta_{2}}\}$. We consider three regions: (a) $\{{\mathbf{X}}_{t}\in D_{1}^{c}\}$, (b) $\{{\mathbf{X}}_{s}\in D_{2}^{c}\}$, and (c) $\{{\mathbf{X}}_{t}\in D_{1},{\mathbf{X}}_{s}\in D_{2}\}$. The large deviation (31) and basic estimates, such as the Potter bound and (6), give a bound of the probabilities of the first two events (a) and (b):
${\mathbb{P}}({\mathbf{X}}_{t}\in D_{1}^{c})+{\mathbb{P}}({\mathbf{X}}_{s}\in D_{2}^{c})\leq\frac{C_{\bf x}}{\ell(t)^{\alpha\theta_{2}-\varepsilon}}.$ (89)
@@
$\begin{array}[]{l}\displaystyle\sum_{{\bf y}\in D_{1}}\sum_{{\bf z}\in D_{2}}{\mathbb{P}}_{\bf x}({\mathbf{X}}_{s}={\bf z},\tau_{{\bf x}}>s)\big|{\mathbb{P}}_{{\bf z}}({\mathbf{X}}_{t-s}={\bf y})-{\mathbb{P}}_{{\bf x}}({\mathbf{X}}_{t-s}={\bf y})\big|\\[5.0pt] \displaystyle\quad\leq\frac{C_{\bf x}B_{s}\hskip 1.0pt\ell(t)^{\theta_{2}}}{B_{t}^{N+1}}\sum_{{\bf y}\in D_{1}}\sum_{{\bf z}\in D_{2}}{\mathbb{P}}_{\bf x}({\mathbf{X}}_{s}={\bf z},\tau_{{\bf x}}>s)\leq\frac{C_{\bf x}B_{s}B_{t}^{N}\hskip 1.0pt\ell(t)^{(N+1)\theta_{2}}}{B_{t}^{N+1}\hskip 1.0ptl(t)^{\upsilon_{N}}},\end{array}$ (90)
where in the last inequality we used that the sum over $\bf z$} is bounded by ${\mathbb{P}}(\tau_{\bf x}>s)\leq C_{\bf x}l(t)^{-\upsilon_{N}}$ thanks to Lemma 3.2. For $\alpha=1$, if we take $\theta_{2}=K\upsilon_{N}+1$ and $\theta_{1}=(N+1)\theta_{2}+1$, then combining the three bounds (87), (89) and (90) we arrive at the desired bound (86). For $\alpha<1$, we take $\theta_{2}$ sufficiently small and $\theta_{1}=(N+1)\alpha\theta_{2}+\alpha$ such that $\theta_{1}<\frac{\alpha}{1-\alpha}$ is satisfied. Then, (87), (89) and (90) imply that there exists some $\delta$ such that (86) holds, which finishes the proof.
Remark 3.4
Via a simple adaption of the proof above, we can change the fixed $\bf x$} to the origin $\bf 0$} such that the same upper bound as in (86) holds for
$\sum_{{\bf y}\in{\mathbb{Z}}^{N}}\Big|{\mathbb{P}}_{\bf x}({\mathbf{X}}_{t}={\bf y},\tau_{{\bf x}}>t)-{\mathbb{P}}_{\bf 0}({\mathbf{X}}_{t}={\bf y}){\mathbb{P}}_{\bf x}(\tau_{{\bf x}}>t)\Big|.$
Indeed, comparing with (57) in Lemma 2.13, we can replace the term ${\mathbb{P}}_{\bf x}({\mathbf{X}}_{t-s}={\bf y})$ in (88) by ${\mathbb{P}}_{\bf 0}({\mathbf{X}}_{t-s}={\bf y})$, and obtain the same upper bound as in (90). Thus, the desired result follows once ${\mathbb{P}}_{\bf x}({\mathbf{X}}_{t}={\bf y})$ and ${\mathbb{P}}_{\bf x}({\mathbf{X}}_{t-s}={\bf y})$ in $R^{\bf y}_{2}$ and $R^{\bf y}_{4}$ are replaced by ${\mathbb{P}}_{\bf 0}({\mathbf{X}}_{t}={\bf y})$ and ${\mathbb{P}}_{\bf 0}({\mathbf{X}}_{t-s}={\bf y})$, respectively.